OpenStudy (anonymous):

What is the length of a the diagonal of a football field that is 100 yards long and 60 yards wide? Explain how you found your answer

OpenStudy (anonymous):

pythagorean theorm

OpenStudy (anonymous):

100^2 +60^2 = c^2 a and b are always the sides that form a right-angle. c is always the hypotenuse. Do you know how to solve this?

OpenStudy (anonymous):

hmmm im not really sure

OpenStudy (anonymous):

10000 + 3600 = c^2 c= sqrt(10000 +3600)

OpenStudy (anonymous):

On that last step, I square-rooted both sides. If you sqrt(c^2), then that makes it c. And then you also have to square-root the right side in order to make things "balanced" or whatever you call it.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

http://www.youtube.com/watch?v=WziqDlzHYxU